OpenStudy (anonymous):

need to find an equation. m=2/3, (5,-2) I got y-(-2)=(2/3)(x-5) and then y+2=2/3x+2 what next?

OpenStudy (anonymous):

shouldn't it be y+2=2x/3-10/3 not y+2=2x/3+2?

OpenStudy (anonymous):

I don't know for sure that is why I am asking.

OpenStudy (anonymous):

ok well when you multiplied it out it was supposed to be what I just gave you. And your next step would then be to subtract 2...6/3 from both sides. That is if your answer should be in y=mx+b format.

OpenStudy (anonymous):

ok i am still confused

OpenStudy (anonymous):

ok what is your confusion?

OpenStudy (anonymous):

well what you are telling me is different than what they are telling me in my book.

OpenStudy (anonymous):

what are they telling you in your book?

OpenStudy (anonymous):

what i just put on here except they are telling is to subtract 1 from each side

OpenStudy (anonymous):

So it's telling you to subtract 1 from both sides as the next step?

OpenStudy (anonymous):

ya and that is why I figured I would ask

OpenStudy (anonymous):

I'm not sure I understand the question because I would have approached it differently to begin with. I would have taken y=mx+b and plugged in your m value of 2/3 and then your point (5, -2) to get -2=2/3(5)+b. Then I would solve for b and get -16/3. Then the equation for the line would be y=(2/3)x-16/3.

OpenStudy (anonymous):

From the info I gave I am suppose to come up with an equation.

OpenStudy (anonymous):

the info you gave being a slope of 2/3 and you're finding the equation of the line passing through the point (5,-2), right?

OpenStudy (anonymous):

ya in the format of y=mx+b

OpenStudy (anonymous):

well I got the answer y=(2/3)x-16/3 with that given information

OpenStudy (anonymous):

and i haven't a clue why the book would tell you to subtract 1 from each side. Are you looking at an example? Where did you get the subtract 1 thing from?

OpenStudy (anonymous):

I just realized how late it is. I have to go to a 6:00 class and I have to leave now to get there on time. I'm really sorry, but I have to go!